# PROTOCOL FOR AN OBSERVATIONAL STUDY ON THE EFFECTS OF COMBINATIONS OF ADVERSE CHILDHOOD EXPERIENCES ON ADULT DEPRESSION

### v1

@@ 5. Design of Data Analysis
Our inference will be based on stratified sets of participants in the BRFSS 2023, as described previously in Section 4. We aim to explore the effect of different combinations of ACEs in our aforementioned candidate sets on depression and also identify those with higher risk. We will deem those combinations of ACEs for which the effect of the treatment is found to be significantly increase the odds of depression in both subgroups - those residing in blue states and red states - as replicable findings. Also, we are interested in those combinations more than doubling the odds ratio of depression in only one subgroup.
@@ 5.1. Data Turnover Framework
We employ data turnover (Bekerman et al. in preparation, Bekerman et al. [2024a]) to analyze the data, ensuring control of the family-wise error rate (FWER; the chance of falsely rejecting any true null hypotheses at level $\alpha=0.05$) and facilitating flexible inference compared to automated cross-screening (Zhao et al. [2018]), by integrating both qualitative and quantitative insights from data exploration and formulating data-informed new hypotheses (e.g., Bekerman et al. [2024b]).
A brief description is as follows. We split our data into two subgroups: those individuals in our study residing in blue states and those residing in red states. The first subgroup, without data exploration, suggests a limited number of combinations of ACEs to be validated in the second subgroup based on some pre-determined selection procedure. Then, we do ad hoc data exploration on the second subgroup to choose how to conduct one-sided hypothesis tests on the first subgroup, possibly allowing for novel hypotheses to be identified and tested. In other words, both subgroups are used to screen and both subgroups are used to validate, but only one subgroup can generate novel hypotheses to be validated in the other subgroup while still maintaining the validity of our entire procedure. Because a limited number of combinations seek validation on the second subgroup, compared with directly controlling the FWER of 385 hypotheses on the whole data (eg. by using Bonferroni correction or Holm procedure), only a small fraction of the 385 potential hypotheses are considered in the validation sample.
@@ 5.2. Analysis Plan
We are interested in testing whether a combination of ACEs would increase the odds of depression compared with the reference population.

### v2

@@ 4. Design of Data Analysis
We aim to explore the effect of different combinations of ACEs on depression and also identify those with higher risk. We will test all of the possible combinations of ACEs, therefore leading to 1024 ($2^{10}$) different hypotheses. We will deem those combinations of ACEs for which the effect more than doubles the odds of depression in both subgroups - those residing in blue states and red states - as replicable findings. Also, we are interested in those combinations more than doubling the odds ratio of depression in only one subgroup, that is testing the global null hypothesis.
@@ 4.1. Data Turnover Framework
We employ data turnover (bekerman2024protocol) to analyze the data, ensuring control of the family-wise error rate (FWER; the chance of falsely rejecting any true null hypotheses at level $\alpha=0.05$) and facilitating flexible inference compared to automated cross-screening (zhao2018cross), by integrating both qualitative and quantitative insights from data exploration and formulating data-informed new hypotheses (e.g., bekerman2024planning).
A brief description is as follows. We split our data into two subgroups: those individuals in our study residing in blue states and those residing in red states. The first subgroup, without data exploration, suggests a limited number of combinations of ACEs to be validated in the second subgroup based on some pre-determined selection procedure. Then, we do ad hoc data exploration on the second subgroup to choose how to conduct one-sided hypothesis tests on the first subgroup, possibly allowing for novel hypotheses to be identified and tested. In other words, both subgroups are used to screen and both subgroups are used to validate, but only one subgroup can generate novel hypotheses to be validated in the other subgroup while still maintaining the validity of our entire procedure. Because a limited number of combinations seek validation on the second subgroup, compared with directly controlling the FWER of 1024 hypotheses on the whole data (eg. by using Bonferroni correction, Holm procedure or directly applying ISS on the whole sample), only a small fraction of the 1024 potential hypotheses are considered in the validation sample.
Moreover, we need to determine whether it is more beneficial to conduct EDA on the smaller or larger sample. Simulation results in Bekerman’s paper (bekerman2024protocol) suggest that performing EDA on the smaller subgroup—in our case, participants living in blue states—is advantageous. We anticipate that incorporating data exploration within this subgroup will strengthen our study’s overall design by enabling data-informed decision-making and facilitating the emergence of new hypotheses.
@@ 4.2. Analysis Plan
As we note before, for each $X_{i}$, we test $H_{0,i}:\eta(X_{i})<\tau\quad\text{versus}\quad H_{1,i}:\eta(X_{i})\geq\tau.$ Therefore, we need to prespecify $\tau$. Intuitively, if $\tau$ is too small, the true superlevel set is very large and its corners tend to lie deep. If $\tau$ is too large, no method has power and all outputs are empty. So we are interested in testing whether a combination of ACEs would more than double the odds of depression compared with the reference population. Let $p_{0}$ be a baseline probability - proportion of people exposed to no ACEs. In previous BRFSS datasets which are not included in this study, $p_{0}$ is around $9.4\%$. To test that the odds at $x$ are at least $c$ times the baseline odds,
$\frac{\tau}{1-\tau}=c\frac{p_{0}}{1-p_{0}}$
$\tau\ =\frac{cp_{0}}{1+(c-1)p_{0}}$
Use $c=2,p_{0}=9.4\%$, $\tau=0.172$.
We plan to integrate isotonic subgroup selection (ISS) with the data turnover method. Recall that in data turnover framework, we split the data into two subgroups $\mathcal{I}_{R}$ (Red Subgroup) and $\mathcal{I}_{B}$ (Blue Subgroup). For every combination $X_{i}$ that we are interested, define the anytime-valid classification $p$-value $p_{\tau}(x)$ as in the binary setting (see Section 3). Below is the pre-specified analysis plan for the red subgroup. Compute red-side $p$-values $p^{R}_{\tau,i}:=p_{\tau}(X_{i};\mathcal{I}_{R})$ for all $i$, and form the screened index set $\mathcal{S}_{\kappa}=\bigl\{\,i:p^{R}_{\tau,i}\leq\kappa\,\bigr\}$ to be tested in $\mathcal{I}_{B}$, for a pre-specified screening cutoff $\kappa\in(0,1)$. From the full DAG $G$, we derive a polyforest $F$ by retaining at most one parent for each node, which can be made at random or based on a distance criterion. We could organize the polyforest $F_{\kappa}^{B}$ in $\mathcal{I}_{B}$ from the information ($p^{R}_{\tau,i}$, $\mathcal{S}_{\kappa}$) we obtain from the screening stage in $\mathcal{I}_{R}$. All nodes of the polyforest $F_{\kappa}^{B}$ are drawn from $\mathcal{S}_{\kappa}$. For each $i\in\mathcal{S}_{\kappa}$, define its set of covers
$\mathrm{Cover}(i)\;=\;\bigl\{\,j\in\mathcal{S}_{\kappa}\,:\,X_{j}\succ X_{i}\ \text{and there is no }k\in\mathcal{S}_{\kappa}\text{ with }X_{j}\succ X_{k}\succ X_{i}\,\bigr\}.$
Select at most one parent for $i$ from $\mathrm{Cover}(i)$ guided by $p$-value: $\mathrm{par}(i)\in\arg\min_{\,j\in\mathrm{Cover}(i)}\ p^{R}_{\tau,j}$, choosing among candidates with tied minimal p-value uniformly at random. If $\mathrm{Cover}(i)=\emptyset$, declare $i$ a root. This yields the polyforest $F_{\kappa}^{B}$ (a DAG with in-degree $\leq 1$). On $F_{\kappa}^{B}$, compute blue-side $p$-values only for screened nodes, $p^{B}_{\tau,i}:=p_{\tau}(X_{i};\mathcal{I}_{B})$ for $i\in\mathcal{S}_{\kappa}$, and run the ISS DAG-testing algorithm with polyforest $F_{\kappa}^{B}$ controlling the FWER for the blue part at level $\alpha/2$.
Using $p_{\tau,i}^{R}$ to choose parents steers each node toward a path whose upstream evidence is strongest, so $\alpha$ is concentrated along promising chains. This mitigates cases where a large-$p$ maximal node would otherwise block the flow of $\alpha$ and prevent strong descendants from ever receiving budget, increasing the power.
Figure 1 illustrates a case where the parent selection creates a bottleneck. Here, nodes $1$, 66, and $7$ are roots, each receiving part of the global $\alpha$-budget (shown in red). Although node $5$ has a small $p$-value ($0.01$), its designated parent is node $6$, whose $p$-value is relatively large ($0.1$). Since node $6$ is unlikely to be rejected with its assigned budget ($0.025$), the rejection cannot propagate to node $5$, and the signal at node $5$ is blocked. This demonstrates how attaching a promising child to an unpromising parent can prevent significant nodes from being discovered.
Figure 1. DAG Testing: Example 1. Each node is labeled with its $p$-value (round brackets); in the induced polyforest-weighted DAG, solid arrows indicate edges that are retained in the polyforest (weight==1), while dashed arrows indicate edges that exist in the full DAG but are not retained in the polyforest (weight==0, since each node can have at most one parent); filled circles indicate hypotheses previously rejected. Panels (a)–(b) correspond to successive iterations of the algorithm, showing how rejections propagate through the structure.
In contrast, Figure 2 shows the situation when node $5$ is attached through node $7$, whose $p$-value is smaller ($0.03$). After node $1$ is rejected in the first iteration ($p_{1}=0.01\leq 0.0125$), the unused budget is reallocated, giving node $7$ a larger share ($0.0333$). With $p_{7}=0.03$, node $7$ can then be rejected, releasing further budget to its child node $5$. Since node $5$ has $p_{5}=0.01$, it is subsequently rejected, and the rejection can continue to its descendants. Simultaneously, since node $6$ is the parent of node $5$ in the full DAG, even if its $p$-value is large, it can also be rejected in line with logical partial order. Compared to Figure 1, the $p$-guided parent selection in Figure 2 avoids the bottleneck at node $6$ and allows the procedure to detect the significant signal at node $5$.
Figure 2. DAG Testing: Example 2. The setting is the same as in Figure 1, except that node $5$ is attached through node 77 rather than node $6$. Panels (a)–(d) correspond to successive iterations of the algorithm, showing how rejections propagate through the structure.